OpenStudy (anonymous):

the sum of twice a number and four less than the number is the differnce -32 and the number. what is the number . shw work

OpenStudy (anonymous):

the second part of problem is not clear, can you rewrite your problem clearly?

OpenStudy (anonymous):

2x+(4-x) = -32-x 2x + 4 -x = -32 - x x + x = -32 - 4 2x = -36 x = -18

OpenStudy (anonymous):

mathmind, you are incorrect. four less than the number means x-4, not 4-x